# The JARVIS Infrastructure is All You Need for Materials Design

### v1

@@ Abstract
Joint Automated Repository for Various Integrated Simulations (JARVIS) is a comprehensive infrastructure offering databases, tools, tutorials, and benchmarks for multiscale, multimodal, forward, and inverse materials design. Emphasizing open access principles and reproducibility, it integrates theoretical and experimental methodologies such as density functional theory, quantum Monte Carlo, tight-binding, classical force fields, and machine-learning approaches-including fingerprinting, graph neural networks, and transformer models. Its experimental data collection spans cryogenics, microscopy, and diffraction, covering materials like metals, semiconductors, insulators, superconductors, carbon capture systems, high-strength compounds, and low-dimensional materials, heterostructures and defects. JARVIS disseminates resources via open datasets, web applications, executable scripts, and peer-reviewed publications, ensuring broad accessibility and reproducibility. Widely adopted worldwide, it has facilitated millions of data and tool downloads. By unifying diverse methods and data under one platform, JARVIS drives both fundamental discoveries and real-world innovations, advancing conventional and data-driven materials design.
[…]
Material Measurement Laboratory, National Institute of Standards and Technology, Gaithersburg, MD 20899, USA
@@ 1 Introduction
Materials design has long been a cornerstone of technological progress, underpinning advancements in energy, computing, healthcare, and manufacturing [1]. Traditionally, materials discovery followed an empirical, trial-and-error approach, often requiring decades to identify and optimize materials for specific applications. The advent of computational materials science, particularly with the rise of density functional theory (DFT), molecular dynamics (MD)/force-field (FF) and machine learning (ML) has significantly accelerated the pace of discovery [2, 3, 4]. However, the growing complexity of materials problems, the vast chemical and structural search space, and the increasing demand for high-performance materials necessitate more sophisticated, integrated frameworks that can streamline the entire process from fundamental property prediction to real-world implementation.
@@
One of the critical challenges in forward materials design is the computational cost associated with methods such as first-principles simulations. While DFT and beyond methods are widely used for predicting electronic and structural properties, it is limited by system size and accuracy trade-offs, particularly for strongly correlated materials and high-temperature phases [14]. In classical approaches, empirical force fields enable large-scale simulations but often lack transferability across different materials and conditions [15]. More recently, machine learning property prediction models and machine learning force fields (MLFFs) have emerged as a promising alternative, offering near-DFT accuracy at significantly lower computational cost, yet challenges remain in ensuring generalizability and robustness [16, 17]. For experimental methods, availability of standardized benchmarks, integration of analytical and computational methods at various lengths and time scales and framework to enhance reproducibility is still a critical challenge [18].
@@
The JARVIS infrastructure [4, 24] distinguishes itself from other materials design frameworks by offering a multimodal, scalable, and reproducible ecosystem that integrates first-principles calculations, machine learning models, experimental datasets, and web-based tools into a unified platform. Databases such as Automatic FLOW for Materials Discovery (AFLOW) [25], Open Quantum Materials Database (OQMD) [10], Novel Materials Discovery (NOMAD) [26], Materials Cloud [27] and open catalyst project (OC20) [28] focus primarily on density functional theory (DFT)-based materials screening. These initiatives have fostered a materials-by-design paradigm, where advanced computation, automation, and artificial intelligence (AI) accelerate the transition from fundamental discoveries to industrial applications for next-generation materials innovation. JARVIS extends its capabilities beyond DFT calculations to include classical force-field, machine learning force fields (MLFF), beyond DFT methods such as quantum Monte Carlo and dynamical mean field theory (DMFT). It also covers all major classes of materials such as superconductors, solar cells, topological materials, low dimensional, catalytic, high-strength, thermoelectric, piezoelectric, dielectric discovery, quantum computing algorithms, and experimental validation datasets, making it a truly multimodal and a multiscale infrastructure. Another key differentiator is reproducibility, which is central to JARVIS through the use of open-access datasets, FAIR-compliant data sharing, and the distribution of workflows via interactive web apps and notebooks, benchmarking platform such as JARVIS-Leaderboard ensuring that users can verify and extend existing research [29]. Additionally, JARVIS is designed for scalability, providing not only extensive high-throughput datasets but also machine learning models capable of accelerating materials discovery at significantly reduced computational cost compared to traditional methods [30].
@@
Unlike various other platforms, which often focus on a single domain (e.g., semiconductors or metals), JARVIS covers a broad range of materials, including superconductors, carbon capture materials, and low-dimensional heterostructures, facilitating both forward and inverse design approaches. Its seamless integration of theoretical and experimental techniques-spanning quantum Monte Carlo, classical force fields, microscopy, and diffraction databases ensures that the infrastructure is not only predictive but also experimentally validated, bridging the gap between computational discoveries and real-world applications. With thousands of users, millions of dataset downloads, and growing adoption across academia, industry, and government research, JARVIS has established itself as a comprehensive infrastructure that accelerates materials innovation through automation, interoperability, and AI-driven insights. In this article, we briefly discuss some of the components, impacts, and outreach of the JARVIS infrastructure. For more detailed information, we refer to our latest references such as [4, 24].
@@ 2.1.1 Databases
JARVIS hosts a suite of materials databases that provide detailed information on various material properties. These databases consist of its own resources such as JARVIS-DFT as well as external datasets such as Alexandria DB [31] in a uniform format to allow seamless integration as well as comparative analysis. The aggregate of such databases can be up to 6 million materials and 10 million properties. These databases have been downloaded close to 2 million times indicating a wide adaptability by the materials research community. Note that there could be overlap among various databases. To name a few of these, the JARVIS-DFT database contains density functional theory calculations for over 80,000 materials, offering data on structural, electronic, optical, mechanical, phonon, and topological properties. Additionally, the JARVIS-FF database includes classical force-field calculations for approximately 2,000 materials, facilitating the study of interatomic interactions and molecular dynamics simulations. These databases are continuously updated to incorporate new materials and properties, ensuring a comprehensive and current resource for researchers. A list of currently available databases are available at: https://pages.nist.gov/jarvis/databases/. While these tutorials are static snapshots of datasets, a much more user friendly and interactive list of databases is available at: https://jarvis.nist.gov/#Databases.
@@ 2.1.2 Tools
To support materials design, JARVIS provides a range of software tools that allow setting up and analysis of materials design tasks, leveraging artificial intelligence (AI) and machine learning (ML) models, as well as simulation frameworks. For instance, JARVIS-Tools allow integrating with more than 15 software types such as Vienna Ab initio Simulation Package (VASP) [32, 33], Quantum Espresso (QE) [34], Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) [15], etc., automatic curation and analysis of results, converting information to webpages, providing links to databases mentioned above etc. Similarly, the Atomistic Line Graph Neural Network (ALIGNN) property predictor [35] utilizes graph neural networks to rapidly predict material properties, while the ALIGNN Force-field tool [36] enables fast structure optimization. All of these tools are open source, available on GitHub and with reasonable documentation to allow both new comers as well as experienced users to utilize resources available in JARVIS. There are hundreds of GitHub stars, and thousands to millions of PyPi downloads of these software tools indicating profound utilization of these novel resources. These tools are designed to streamline the materials design process, thereby reducing the reliance on time-consuming methods such as first-principles calculations. Currently, there are more than 10 such tools available through JARVIS. A collection of such tools can be found at the https://github.com/knc6
@@ 2.1.3 Tutorials
Recognizing the importance of education and accessibility, JARVIS offers a collection of Jupyter, Google Colab and Serverless Materials Design (SLMat) [37] notebooks that serve as tutorials for various methods in materials design. These resources cover topics such as electronic structure calculations, force-field development, AI applications, and quantum computing methods. By providing these educational materials, JARVIS aims to equip researchers and students with the necessary skills to effectively utilize its tools and databases in their work. One of the major components of tutorials is the JARVIS-Tools-Notebooks framework https://github.com/JARVIS-Materials-Design/jarvis-tools-notebooks available on GitHub. Currently, there are more than 100 notebooks available in this repository. Another set of documented tutorials are available at https://pages.nist.gov/jarvis/tutorials/ .
@@ 2.1.4 Benchmarking Standards
To ensure the reliability and reproducibility of materials design methods, JARVIS has developed the JARVIS-Leaderboard https://pages.nist.gov/jarvis_leaderboard/, an open-source, community-driven platform that facilitates benchmarking across various categories, including AI, electronic structure, force-fields, quantum computation, and experimental methods. This platform allows users to set up custom benchmarks and contribute datasets, code, and metadata, promoting transparency and standardization in materials research [29]. Currently, there are more than 8 million data points and 1300 benchmarks in the leaderboard and it is continuously expanding.
@@ 3 Theoretical and Experimental Method Integration
Figure 2: Evolution of Different Projects in the NIST JARVIS Infrastructure (2017–2025). This figure illustrates the timeline and development of various JARVIS initiatives, including JARVIS-DFT, JARVIS-FF, JARVIS-ML, JARVIS-Exp, and other related projects within the NIST JARVIS ecosystem. The steady expansion from 2017 to 2025 highlights the integration of density functional theory (DFT) databases, machine learning models, force field developments, and experimental data repositories. The evolution reflects increasing computational capabilities, data-driven approaches, and interdisciplinary collaborations aimed at advancing materials design and discovery.
@@ 3.1.1 Density Functional Theory and Tight-binding
Because of its ab-inito nature and wide usability, DFT is a cornerstone of JARVIS. Density Functional Theory (DFT) is a computational quantum mechanical modeling method employed to investigate the electronic structure of many-body systems. The JARVIS-DFT database encompasses a vast array of materials, offering detailed information on structural, electronic, optical, and mechanical properties [39, 40, 41, 4], thereby serving as a valuable resource for researchers in materials science. The DFT databases in JARVIS are JARVIS-DFT which is mainly based on VASP calculations and JARVIS-QETB databases which is based on Quantum Espresso (QE) based calculations. DFT has various levels of fidelity such as local-density approximations (LDA), generalized gradient approximation (GGA), metaGGA, van der Waals (vdW) corrected functionals such as OptB88vdW and many body based methods which have various levels of computational cost. JARVIS encompasses most of these techniques in JARVIS-DFT, JARVIS-QETB as well in the JARVIS-Leaderboard framework. JARVIS-QETB was used to develop a universal set of tight-binding parameters across the periodic table [42]. It includes both two-body and three-body effective interaction terms, plus self-consistent charge transfer, enabling the model to work for metallic, covalent, and ionic bonds with the same parameter set.

### v2

@@ Abstract
The Joint Automated Repository for Various Integrated Simulations (JARVIS) is a unified platform for multiscale, multimodal, forward, and inverse materials design. It integrates diverse theoretical and experimental approaches, including density functional theory, quantum Monte Carlo, tight-binding, classical force fields, machine learning, microscopy, diffraction, and cryogenics, across a wide range of materials. Emphasizing open access and reproducibility, JARVIS provides datasets, tools, benchmarks, and web applications that are widely adopted by the materials community. By bridging computation and experiment, JARVIS accelerates both fundamental research and real-world materials innovation.
@@ 1 Introduction
Materials design has long been a cornerstone of technological progress, underpinning advancements in energy, computing, healthcare, and manufacturing [1, 2, 3]. Traditionally, materials discovery followed an empirical, trial-and-error approach, often requiring decades to identify and optimize materials for specific applications. The advent of computational materials science, particularly with the rise of density functional theory (DFT), tight-binding (TB), molecular dynamics (MD)/force-field (FF) and machine learning (ML) has significantly accelerated the pace of discovery [4, 5, 6]. However, the growing complexity of materials problems, the vast chemical and structural search space, and the increasing demand for high-performance materials necessitate more sophisticated, integrated frameworks that can streamline the entire process from fundamental property prediction to real-world implementation.
@@
One of the critical challenges in forward materials design is the computational cost associated with methods such as first-principles simulations. While DFT and beyond methods are widely used for predicting electronic and structural properties, they are limited by system size and accuracy trade-offs, particularly for strongly correlated materials and high-temperature phases [16]. In classical approaches, empirical force fields enable large-scale simulations but often lack transferability across different materials and conditions [17]. More recently, machine learning property prediction models and machine learning force fields (MLFFs) have emerged as a promising alternative, offering near-DFT accuracy at significantly lower computational cost, yet challenges remain in ensuring generalizability and robustness [18, 19]. For experimental methods, availability of standardized benchmarks, integration of analytical and computational methods at various lengths and time scales and framework to enhance reproducibility is still a critical challenge [20].
@@
The JARVIS infrastructure [6, 26] distinguishes itself from other materials design platforms by offering a multimodal, multiscale, and reproducible ecosystem that integrates first-principles calculations, machine learning models, experimental datasets, and interactive web-based tools into a unified framework.
Prominent materials informatics resources such as the Automatic FLOW for Materials Discovery (AFLOW) [27], Materials Project [4], Open Quantum Materials Database (OQMD) [12], Novel Materials Discovery (NOMAD) [28], Materials Cloud [29], and the Open Catalyst Project (OC20) [30] have significantly advanced DFT-based materials screening. However, they typically focus on a single modeling scale and do not broadly incorporate methods such as tight-binding, classical force-fields, dynamical mean field theory (DMFT), quantum Monte Carlo (QMC), microscopy, cryogenic measurements, and quantum computing techniques. In contrast, JARVIS integrates these diverse modeling strategies to support a true multiscale design workflow. For example, JARVIS-QETB supports tight-binding parameterization; JARVIS-FF and ALIGNN-FF enable large-scale simulations with classical and machine-learned force fields; and JARVIS-QC explores quantum computation algorithms for materials modeling. Furthermore, JARVIS encompasses a wide range of material classes and applications, including superconductors, solar cells, thermoelectrics, piezoelectrics, dielectrics, low-dimensional materials, catalytic systems, topological insulators, and datasets for experimental validation.
A key differentiator is reproducibility-JARVIS provides open-access, FAIR-compliant datasets and workflows distributed via web applications, notebooks, and the JARVIS-Leaderboard [31], which facilitates rigorous benchmarking across multiple domains. This ensures that results can be independently verified and extended by the community. Finally, JARVIS is designed for scalability, offering not only high-throughput DFT and force-field data, but also AI-driven models (e.g., ALIGNN, AtomGPT) that can accelerate materials prediction at significantly reduced computational cost compared to conventional approaches [32].
@@
Unlike various other platforms, which often focus on a single domain such as semiconductors, metals, or alloys, JARVIS encompasses a broad range of materials, including superconductors, carbon capture frameworks, and low-dimensional heterostructures. It supports both forward and inverse design approaches and integrates a diverse set of theoretical and experimental techniques-spanning quantum Monte Carlo, classical force fields, microscopy, and diffraction data to ensure that predictions are not only computationally robust but also experimentally validated. This multimodal integration bridges the gap between in silico discoveries and real-world applications.
While the prediction and identification of new compounds are central to JARVIS, the infrastructure also facilitates the full materials design process. It enables multiscale property prediction across quantum, classical, and experimental domains; supports targeted materials optimization using models such as ALIGNN and AtomGPT, often guided by leaderboard-based performance rankings; and promotes reproducibility and transparency through rigorous benchmarking protocols. Workflow automation, enabled by the JARVIS-tools Python package, ensures consistent and reproducible high-throughput simulations, while FAIR-compliant data practices promote interoperability and reuse. The infrastructure also incorporates rich, multimodal datasets, including those derived from microscopy, spectroscopy, and diffraction, providing a holistic environment for materials research.
With thousands of users, millions of dataset downloads, and expanding adoption in academic, industrial, and governmental settings, JARVIS has emerged as a comprehensive infrastructure that accelerates materials innovation through automation, openness, and AI-driven insights. This article provides an overview of selected components, impacts, and community engagement within the JARVIS ecosystem. For further details, we refer readers to our prior publications [6, 26].
@@ 2.1.1 Databases
JARVIS hosts a suite of materials databases that provide detailed information on various material properties. These databases consist of its own resources such as JARVIS-DFT as well as external datasets such as Alexandria DB [33] in a uniform format to allow seamless integration as well as comparative analysis. The aggregate of such databases can be up to 6 million materials and 10 million properties. These databases have been downloaded close to 2 million times indicating a wide adaptability by the materials research community. Note that there could be overlap among various databases. To name a few of these, the JARVIS-DFT database contains density functional theory calculations for over 90,000 materials, offering data on structural, electronic, optical, mechanical, phonon, and topological properties. Additionally, the JARVIS-FF database includes classical force-field calculations for approximately 2,000 materials, facilitating the study of interatomic interactions and molecular dynamics simulations. These databases are continuously updated to incorporate new materials and properties, ensuring a comprehensive and current resource for researchers. A list of databases are available at: https://atomgptlab.github.io/jarvis-tools/databases/. While these tutorials are static snapshots of datasets, a much more user friendly and interactive list of databases is available at: https://jarvis.nist.gov/#Databases.
@@ 2.1.2 Tools
To support materials design, JARVIS provides a range of software tools that allow setting up and analysis of materials design tasks, leveraging artificial intelligence (AI) and machine learning (ML) models, as well as simulation frameworks. For instance, JARVIS-Tools allow integrating with more than 15 software types such as Vienna Ab initio Simulation Package (VASP) [34, 35], Quantum Espresso (QE) [36], Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) [17], etc., automatic curation and analysis of results, converting information to webpages, providing links to databases mentioned above etc. Similarly, the Atomistic Line Graph Neural Network (ALIGNN) property predictor [37] utilizes graph neural networks to rapidly predict material properties, while the ALIGNN Force-field tool [38] enables fast structure optimization. All of these tools are open source, available on GitHub and with reasonable documentation to allow both new comers as well as experienced users to utilize resources available in JARVIS. There are hundreds of GitHub stars, and thousands or millions of PyPi downloads of these software tools indicating profound utilization of these novel resources. These tools are designed to streamline the materials design process, thereby reducing the reliance on time-consuming methods such as first-principles calculations. Currently, there are at least 13 such tools available through JARVIS as shown in Table. 1.
Table 1: JARVIS-related repositories and toolkits for materials design under the GitHub organization: https://github.com/atomgptlab .
Name Short Description
1.jarvis-tools JARVIS-Tools: An open-source software package for data-driven atomistic materials design
2.alignn ALIGNN: Atomistic Line Graph Neural Network and force-field
3.atomgpt AtomGPT: Atomistic Generative Pretrained Transformer for Forward and Inverse Materials Design
4.chemnlp ChemNLP: A Natural Language Processing based Library for Materials Chemistry Text Data
5.atomvision AtomVision: Deep learning framework for atomistic image data
6.atomqc AtomQC: Atomistic Calculations on Quantum Computers
7.tb3py TB3Py: Two- and three-body tight-binding calculations for materials
8.intermat InterMat: Interface materials design toolkit
9.defectmat DefectMat: Defect materials design toolkit
10.catalysismat CatalysisMat: Catalytic materials design toolkit
11.jarvis-tools-notebooks A Google-Colab Notebook Collection for Materials Design
12.jarvis_leaderboard JARVIS-Leaderboard: Explore State-of-the-Art Materials Design Methods and Reproducible Benchmarks
@@ 2.1.3 Tutorials
Recognizing the importance of education and accessibility, JARVIS offers a collection of Jupyter, Google Colab and Serverless Materials Design (SLMat) [39] notebooks that serve as tutorials for various methods in materials design. These resources cover topics such as electronic structure calculations, force-field development, AI applications, and quantum computing methods. By providing these educational materials, JARVIS aims to equip researchers and students with the necessary skills to effectively utilize its tools and databases in their work. One of the major components of tutorials is the JARVIS-Tools-Notebooks framework https://github.com/atomgptlab/jarvis-tools-notebooks available on GitHub. Currently, there are more than 100 notebooks available in this repository. Another set of documented tutorials are available at https://atomgptlab.github.io/jarvis-tools/tutorials/ .
@@ 2.1.4 Benchmarking Standards
Rigorous, transparent benchmarking is essential for reproducible materials research. To this end, JARVIS hosts the JARVIS‑Leaderboard (https://pages.nist.gov/jarvis_leaderboard/), an open-source, community‑driven platform that evaluates methods spanning the full materials‐design workflow [31]. The leaderboard is built with mkdocs, version‑controlled on GitHub, and rebuilt automatically via continuous‑integration (CI) tests, ensuring that every contribution can be reproduced from code, data and metadata.
At the time of writing, the leaderboard contains 322 benchmarks, 2087 contributions, and more than 8.75 million individual data points drawn from 503 distinct methods and maintained by 26 active contributors. Benchmarks are organised into five top‑level categories-Artificial Intelligence (AI), Electronic Structure (ES), Force-Fields (FF), Quantum Computation (QC) and Experiments (EXP) with contributions currently distributed in following numbers: AI 1034, ES 741, FF 282, QC 6 and EXP 25.
Each task is defined by a machine‑readable folder that bundles: (i) a ground‑truth dataset (JSON/CSV, optionally zipped), (ii) one or more evaluation metrics (e.g. MAE, RMSE, accuracy, MULTIMAE/L1 norm for spectra), and (iii) minimal metadata (metadata.json) describing the target property, units, data split, licence and citation. The detailed contributor guide (https://pages.nist.gov/jarvis_leaderboard/guide/guide/) walks users through forking the GitHub repository, formatting predictions as *.csv.zip, adding a run.sh/Dockerfile for full reproducibility, and submitting a pull‑request that triggers CI tests and eventually publishes results on the static. At present, the primary goal of the JARVIS-Leaderboard is to provide a consistent and reproducible benchmarking framework for comparing models and methods across multiple material properties. While it does not currently track historical changes in property predictions for individual compounds, this is a valuable direction for future development.
@@ 3 Theoretical and Experimental Method Integration
Figure 2: Evolution of Different Projects in the NIST JARVIS Infrastructure (2017-2025). This figure illustrates the timeline and development of various JARVIS initiatives, including JARVIS-DFT, JARVIS-FF, JARVIS-ML, JARVIS-Exp, and other related projects within the NIST JARVIS ecosystem. The steady expansion from 2017 to 2025 highlights the integration of density functional theory (DFT) databases, machine learning models, force field developments, and experimental data repositories. The evolution reflects increasing computational capabilities, data-driven approaches, and interdisciplinary collaborations aimed at advancing materials design and discovery.
@@ 3.1 Quantum Mechanical Methods
JARVIS-DFT MP OQMD
Materials 90000 144595 1022663
DFT methods OptB88vdW,mBJ,SOC PBE,SCAN PBE
K-point/cut-off Converged per mat. Fixed Fixed
SCF convergence criteria Energy & Forces Energy Energy
Elastic tensors/phonons 17402 14072 -
Piezoelectric, IR spect. 4801 3402 -
Dielectric tensors (w/o ion) 4801 (15860) 3402 -
Electric field gradients 11865 - -
XANES spectra - 22000 -
2D monolayers 1011 - -
Raman spectra 400 50 -
Seebeck, Power F 23210 48000 -
Solar SLME 8614 - -
SOC Spillage 11383 - -
WannierTB 1771 - -
STM images 1432 - -
Supercon Tc 2200 - -
Vacancy 400/192494 - -
Surfaces 300 - -
Interfaces 600/1.4 trill. - -
Table 2: Comparison of three major computational DFT-based materials databases: JARVIS-DFT, Materials Project (MP), and the Open Quantum Materials Database (OQMD). The table summarizes key characteristics and property coverage across the platforms, highlighting both general metrics (e.g., number of materials and DFT methods) and specific physical properties computed.
@@ 3.1.1 Density Functional Theory and Tight-binding
Because of its ab initio nature and wide applicability, Density Functional Theory (DFT) serves as a cornerstone of the JARVIS infrastructure. DFT is a quantum mechanical modeling method used to investigate the electronic structure of many-body systems, particularly in condensed matter physics and materials science. The JARVIS-DFT database provides a comprehensive repository of computed structural, electronic, optical, and mechanical properties for a wide array of materials [41, 42, 43, 6], making it a valuable resource for the materials research community.
JARVIS includes two primary DFT-based datasets: JARVIS-DFT, primarily built using the Vienna Ab-initio Simulation Package (VASP), and JARVIS-QETB, based on Quantum ESPRESSO calculations. These databases span a range of DFT fidelities, including local density approximation (LDA), generalized gradient approximation (GGA), meta-GGA, van der Waals (vdW)-corrected functionals such as OptB88vdW, and beyond, including many-body perturbation theory methods. The JARVIS-Leaderboard framework further standardizes these calculations for benchmarking and model validation.
The JARVIS-QETB database was also instrumental in the development of a universal tight-binding model across the periodic table [44]. This model includes two-body and three-body effective interaction terms along with self-consistent charge transfer, allowing it to accurately describe metallic, covalent, and ionic bonds with a unified parameter set.
To contextualize the role of JARVIS among other leading DFT databases, Table 2 provides a comparison of three prominent platforms: JARVIS-DFT, the Materials Project (MP), and the Open Quantum Materials Database (OQMD). The table summarizes key characteristics and the breadth of physical properties computed across platforms. JARVIS-DFT emphasizes per-material convergence of k-point meshes and plane-wave cutoffs, and includes a number of uniquely reported quantities such as electric field gradients, spin-orbit coupling (SOC) spillage, two-dimensional monolayers, scanning tunneling microscopy (STM) images, and superconducting transition temperatures. In contrast, MP features a broad range of high-throughput properties including XANES spectra and thermoelectric transport coefficients, while OQMD prioritizes scale, offering formation energies for over one million materials. Numbers in parentheses in the table indicate data overlaps, such as the 41,697 structures common to both JARVIS and MP. The vacancy and interface rows denote the number of computed entries and machine learning model based predictions (e.g., 400 vacancy defects with DFT and 192,494 ALIGNN based predictions; 600 computed interfaces with alternating slab junction model using DFT and 1.4 trillion interfaces analyzed with ALIGNN). Together, these databases exhibit complementary strengths: JARVIS prioritizes depth and diversity of physical property predictions, MP provides extensive high-throughput coverage, and OQMD offers a broad-scale energetics dataset. Note that all platforms are continuously evolving, and database sizes are subject to change.